OpenStudy (anonymous):

I cam across this sentence that I need help understanding: "Since G is abelian (communatative), (ab^-1)=ab^-1*ab^-1"

OpenStudy (anonymous):

What dose being communative have to do with (ab^-1)=ab^-1*ab^-1"

OpenStudy (kmeis002):

I'm assuming ab is denoted the operation on two unique elements a and b and then taking the inverse after this operation: \[(a*b)^{-1} \] It this correct?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

I know communtaivety is ab=ba, but what does that have to do with (ab^-1)=ab^-1*ab^-1"

OpenStudy (anonymous):

(ab^-1)=ab^-1*ab^-1=a^2(b^-1)^2=a^2(b^2)^-1=ee^-1=e

OpenStudy (anonymous):

that was the rest of it

OpenStudy (kmeis002):

Does the text of the book look like this (the parenthesis is throwing me off) \[ (ab)^{-1} = ab^{-1}*ab^{-1} \]

OpenStudy (anonymous):